### v1

# What is and is not inside a Cayley graph?

Abstract
In this note we show that there is a cubic graph of girth $5$ that is not a subgraph of any minimal Cayley graph. On the other hand, we show that any Generalized Petersen Graph $G(n,k)$ with $\gcd(n,k)=1$ is an induced subgraph of a minimal Cayley graph. These results give insights into two comments of László Babai in [L. Babai, Automorphism groups, isomorphism, reconstruction. Graham, R. L. (ed.) et al., Handbook of combinatorics. Vol. 1-2, 1994].
1 Introduction
Given a finite group $\Gamma$ and a connection set $S\subseteq\Gamma$, the (undirected, right) Cayley graph $\textrm{Cay}(\Gamma,S)$ has vertex set $\Gamma$, where two elements $a,b\in\Gamma$ form an edge if $a^{-1}b\in S$. Note that we consider Cayley graphs as undirected even if the connection set is not inverse-closed.
A Cayley graph is minimal (also known as irreducible) if $S$ is an inclusion-minimal generating set of $\Gamma$. Minimal Cayley graphs arise naturally: a famous and open problem, often attributed to Lovász, is (equivalent to) the question of whether minimal Cayley graphs are Hamiltonian; see [21, Section 4]. Other contexts in which minimal Cayley graphs appear include the study of the genus of a group [28] and problems related to sensitivity [10, Questions 7.7 and 8.2]. Various aspects of these graphs—sometimes for specific groups—and their distinguishing features compared to general Cayley graphs have been explored in [19, 23, 17, 15, 13, 8].
A major question, raised by Babai [1, 3], with only partial recent results [12, 7] is whether minimal Cayley graphs have bounded chromatic number. One way to construct minimal Cayley graphs of large chromatic number, would be to take a known family of graphs of large chromatic number and show that its members are subgraphs of minimal Cayley graphs. However, while every graph is an induced subgraph of some Cayley graph [4, 14, 2], this is far from true for minimal Cayley graphs. For instance, neither the complete graph on four vertices minus an edge, $K_{4}-e$, nor the complete bipartite graph $K_{3,5}$ can be subgraphs of a minimal Cayley graph [1]. Moreover, for every $g\geq 3$, there exists a graph of girth $g$ and maximum degree $100$ that is not a subgraph of any minimal Cayley graph [25]. A natural question is how much the degree in the above result can be lowered. We quote from [3]:
It is not known whether or not such excluded subgraphs of girth 5 and degree 3 exist […] The Petersen graph is a subgraph of a minimal Cayley graph of a group of order 20.
In this note we give an answer to the first part of the above statement, i.e., we provide a graph of girth 5 and degree 3 that is not a subgraph of a minimal Cayley graph. Further we extend the second part of the above statement, by showing that any Generalized Petersen Graph $G(n,k)$ with $\gcd(n,k)=1$ is an induced subgraph of a minimal Cayley graph.
2 No-lonely-color colorings of digraphs
It was known since [1] that subgraphs of minimal Cayley graphs admit so-called no-lonely-color colorings of their edges. However, this necessary property was not sufficient to find a cubic girth $5$ graph, that is not a subgraph of a Cayley graph. Hence, we need a stronger necessary condition. This is what we provide in this section.
For a group $\Gamma$ and subset $S$ the Cayley digraph $\overrightarrow{\textrm{Cay}}(\Gamma,S)$ has vertex set $\Gamma$, where two elements $a,b\in\Gamma$ form an arc $(a,b)$ if $a^{-1}b\in S$. Note that antiparallell arcs, i.e., of the form $(a,b)$ and $(b,a)$, may occur. We say that $D$ is a minimal Cayley digraph if there is an inclusion-minimal generating set if sone group $\Gamma$ such that $D\cong\overrightarrow{\textrm{Cay}}(\Gamma,S)$. Cleary,
Observation 2.1.
A digraph $D$ is subdigraph of a minimal Cayley digraph if and only if the underlying graph $G$ of $D$ is subgraph of a minimal Cayley graph.
Denote by $W=(a^{\sigma(1)}_{1},\ldots,a^{\sigma(\ell)}_{\ell})$ a walk in $D$, where the sign $\sigma(i)\in\{+1,-1\}$ denotes wether arc $a_{i}$ is traversed in forward or backward direction respectively. Finally, denote by $\underline{W}$ the subdigraph of $D$ on the set of arcs $\{a_{1},\ldots,a_{\ell}\}$.
Definition 2.1.
An arc-coloring $f:A\to S$ of a digraph $D=(V,A)$ is called faithful no-lonely-color if
  1. 1.
    for every $i\in S$ the subdigraph $f^{-1}(i)$ has maximum outdegree and maximum indegree at most $1$
  2. 2.
    for every cycle $C\subseteq A$ of $D$ and $i\in S$ we have $|C\cap f^{-1}(i)|\neq 1$
  3. 3.
    for walks $W=(a^{\sigma(1)}_{1},\ldots,a^{\sigma(\ell)}_{\ell}),W^{\prime}=(b^{\tau(1)}_{1},\ldots,b^{\tau(\ell)}_{\ell})$ we have $f(a_{i})=f(b_{i})$ and $\sigma(i)=\tau(i)$ for every $i\in[\ell]$, then $\underline{W}\cong\underline{W^{\prime}}$.
Observation 2.2.
Let $f:A\to S$ be a faithful no-lonely-coloring of $D=(V,A)$ then for any $i\in S$
  1. 1.
    reversing the orientation of $f^{-1}(i)$ yields a digraph $D^{i}$ for which $f$ also is a faithful no-lonely-coloring.
  2. 2.
    all cycles in $f^{-1}(i)$ are directed (by Definition 2.1.1.) and monchromatic (by Definition 2.1.3.).
Lemma 2.3.
If $D=(V,A)$ is subdigraph of a minimal Cayley digraph $\overrightarrow{\textrm{Cay}}(\Gamma,S)$, then assigning via $f:A\to S$ to each $a\in A$ the corresponding element of $S$ is a faithful no-lonely-color.
Proof.
We show the properties one by one: 1. For every $i\in S$ the subdigraph $f^{-1}(i)$ consists of the edges corresponding to one element of $c\in S$. Hence, every vertex $g$ can have at most one outneighbor $vc$ and one in-neighbor $vc^{-1}$.
2. By Observation 2.1 the underlying graph of $D$ is subgraph of a minimal Cayley graph. By Babai [1] the coloring $f$ satisfies 2.
3. Consider for walks $W=(a^{\sigma(1)}_{1},\ldots,a^{\sigma(\ell)}_{\ell}),W^{\prime}=(b^{\tau(1)}_{1},\ldots,b^{\tau(\ell)}_{\ell})$ we have $f(a_{i})=f(b_{i})$ and $\sigma(i)=\tau(i)$ for every $i\in[\ell]$. Now let $g$ and $h$ be the starting vertices of $W$ and $W^{\prime}$, respectively. It is well-known and easy to see that in $\overrightarrow{\textrm{Cay}}(\Gamma,S)$ the left-multiplication with $hg^{-1}$ is a digraph automorphism that preserves arc-colors and in particular maps $g$ to $h$, see e.g. [16, Chapter 7.3]. Since $f(a_{i})=f(b_{i})$ and $\sigma(i)=\tau(i)$ for every $i\in[\ell]$, this automorphism induces an isomorphism from $\underline{W}$ to $\underline{W^{\prime}}$. ∎
It follows directly that
Observation 2.4.
Let $D=(V,A)$ be a subgraph of a minimal Cayley digraph $\overrightarrow{\textrm{Cay}}(\Gamma,S)$ and $f:A\to S$ the faithful no-lonely-color coloring coming from Lemma 2.3. If for some $i\in S$ there is a cycle of length $2$ in $f^{-1}(i)$, then each component of $f^{-1}(i)$ has at most two vertices and augmenting it to a cycle of length $2$, produces a $D^{\prime}$, such that $D\subseteq D^{\prime}\subseteq\overrightarrow{\textrm{Cay}}(\Gamma,S)$ with the correspondingly extended $f$.
We turn our attention to undirected graphs.
Definition 2.2.
An edge-coloring $f:E\to S$ of a graph $G=(V,E)$ is called faithful no-lonely-color if
  1. 1.
    for every $i\in S$ the subgraph $f^{-1}(i)$ has maximum degree $2$,
  2. 2.
    for every cycle $C\subseteq E$ of $G$ and $i\in S$ we have $|C\cap f^{-1}(i)|\neq 1$,
  3. 3.
    for every $i\in S$ all cycles in the subgraph $f^{-1}(i)$ are of the same length.
If $G$ is subgraph of a minimal Cayley graph, then by Observation 2.1 $G$ is the underlying graph of $D$ a subdigraph of a minimal Cayley digraph. Now, we can apply Lemma 2.3 to $D$. By Observatiobn 2.2.2 we get:
Lemma 2.5.
If $G=(V,E)$ is subgraph of a minimal Cayley graph $\textrm{Cay}(\Gamma,S)$, then assigning via $f:E\to S$ to each $e\in E$ the corresponding element of $S$ yields a faithful no-lonely-color.
A useful reformulation of Definition 2.2.2 is the following:
Observation 2.6.
Let $f:E\to S$ be an edge-coloring of $G=(V,E)$. Then $f$ satisfies Definition 2.2.2 if and only if for any $i\in S$ and $e=\{u,v\}\in f^{-1}(i)$ the vertices $u,v$ lie in different connected components of $G\setminus f^{-1}(i)$.
3 An algorithm
The implementation of the following algorithms, is available in our GitHub repository [24].
First, we devise an algorithm that given a graph $G=(V,E)$ produces all faithful no-lonely-color colorings. We extend a partial coloring $f$. We assume the set $C=\{1,\ldots,k\}$ for some $k\in\mathbb{N}$. Initially however we set, $f(e)=0$ for all $e\in E$, which stands for $e$ being uncolored.
1
2Input:
[…]
5
6if all edges are colored and $f\_valid$ then
7       Output current coloring $f$
8       return True
// Determine candidate edges to color next
9 if there exists a cycle with a unique uncolored edge $e$ then
10       $next\_edges\leftarrow\{e\}$
11 else
12       if $f\_valid$ then
             // Start coloring with a new color $c+1$
13             Choose any uncolored edge $e$
14             $f(e)\leftarrow c+1$
15             if extendColoring($c+1$, False) then
16                   return True
            // Try other uncolored edges with current color
17             $next\_edges\leftarrow$ uncolored edges of $G$
18       else
             // Fix a cycle with exactly one edge of color $c$
19             Find a cycle $C$ with a unique edge of color $c$
20             if such a cycle $C$ exists then
21                   $next\_edges\leftarrow$ uncolored edges of $C$
22             else
23                   $next\_edges\leftarrow$ uncolored edges of $G$
24foreach $e\in next\_edges$ do
25       $f(e)\leftarrow c$
26
27      if $\Delta(f^{-1}(c))\leq 2$ and all cycles in $f^{-1}(c)$ have the same length then
             // Check Observation 2.6
28             if $u^{\prime},v^{\prime}$ are in different components of $G\setminus f^{-1}(c)$ for all $e^{\prime}\in f^{-1}(c)$ then
29                   if extendColoring($c$, True) then
30                         return True
31            if extendColoring($c$, False) then
32                   return True
33      else
             $f(e)\leftarrow 0$ ;
              // Reset edge if invalid
34
return False
Algorithm 1 extendColoring: Extends a partial faithful no-lonely-color coloring
After having generated all faithful no-lonely-color colorings of $G$ we reduce the list by removing colorings that can be obtained from others by applying some automorphism of $G$ or by relabeling colors. We then proceed to find orientations of $G$ that satisfy Lemma 2.3 and in particular Observation 2.4.
For the following assume we are given $G=(V,E)$ together with a no-lonely-color coloring $f:E\to S$. We devise an algorithm to construct an orientation $D$ of $G$ color by color to obtain a candidate orientation. It is based on the following easy observations. For this, we note that if $H$ is connected and non-trivial, i.e., with more than one vertex of maximum degree $\leq 2$. Then $H$ has exactly two orientations satisfying Definition 2.1.1, denoted by $H^{+}$ and $H^{-}$. By Observation 2.2 in each color, one of these orientations can be arbitrarily fixed. Further, note that by Obervation 2.4, only if some color $f^{-1}(i)$ is a matching, then its edges can potentially be completely bi-oriented. We initialize $D=G$ unoriented and $c=1$.
1 Input: An integer $c$ indicating the next color class to orient.
2
3if $c=k+1$ then
4       output $D$
5else
6       fix an arbitrary non-trivial connected component $H^{\prime}$ of $f^{-1}(c)$
7       for $H^{\prime}\in S\subseteq\{H\mid H\text{ is a non-trivial connected component of }f^{-1}(c)\}$ do
8             Orient each $H\in S$ as $H^{+}$ in $D$
9             Orient each $H\notin S$ as $H^{-}$ in $D$
10             extendCandidateOrientation$(c+1)$
11      if $f^{-1}(c)$ is a matching then
12             Bi-orient all edges in $f^{-1}(c)$ in $D$
13             extendCandidateOrientation$(c+1)$
Algorithm 2 extendCandidateOrientation: Extends candidate orientations of a graph with no-lonely-color $k$-coloring.
For each of the generated candidate orientations we use the following algorithm that checks the special case of Definition 2.1.3, where only cycles are considered.
1 Input:
[…]
3 Output:
4True if, for every cycle $C$ in $D$, the color/direction pattern of $C$ starting at any vertex yields a cycle; False otherwise.
5for every cycle $C$ in $D$ do
6       Represent $C$ as a sequence of arcs with directions:
7        $C=(a^{\sigma(1)}_{1},\ldots,a^{\sigma(\ell)}_{\ell})$
8       where each $a_{i}$ is an arc and $\sigma(i)\in\{+1,-1\}$ indicates direction. Let $\mathbf{c}=(f(a_{1}),\ldots,f(a_{\ell}))$ be the color sequence.
9       Let $\boldsymbol{\sigma}=(\sigma(1),\ldots,\sigma(\ell))$ be the direction sequence. for every vertex $v$ in $V$ do
10             Try to follow a path from $v$ using color sequence $\mathbf{c}$ and direction sequence $\boldsymbol{\sigma}$.
11             if a path is found, but it does not return to $v$ or self-intersects then
12                   return False
return True
Algorithm 3 checkCycles: Verify if the color sequence of every cycle, when followed from any vertex, yields a cycle
4 Two cubic graphs of grith 5
There are two cubic graphs of grith 5. Our implementation [24] applied to the triplex graph, see the left of Figure 1, shows that it is not the subgraph of a minimal Cayely graph. Indeed, Algorithm 1 produces all 216 faithful no-lonely color colorings associated to this graph, in 11 seconds. After reducing modulo automorphism, we are only left with 15 colorings. Algorithm 2 and Algorithm 3 finish without finding any orientation to any of these colorings in about 45 seconds.
Figure 1: Left: The triplex graph is not subgraph of a minimal Cayley graph. Right: An orientation of the twinplex graph with a coloring from our algorithm.
The second cubic graphs of grith 5 is the twinplex graph depicted on the right of Figure 1. In this case, our procedure finds an orientation and a coloring that passes our algorithm, as depicted in the figure. Hence, we cannot conclude whether the graph is a subgraph of a minimal Cayley graph or not.
5 Generalized Petersen Graphs
[…]
Figure 2: Left: the filled vertices induce $G(7,2)$ in $\textrm{Cay}(C_{7}\rtimes C_{3},\{a,b\})$. Right: The Petersen plane, where white dots are Cayley graphs, Grey dots are induced subgraph of a minimal Cayley graph of $C_{n}\rtimes C_{r}$ by Theorem 5.1, and black dots have been obtained by computer.
Theorem 5.1.
Let $k,r,n\in\mathbb{N}$ such that $\gcd(k,n)=1$ and $k^{r}=1\mod n$. Then $G(n,k)$ is an induced subgraph of a minimal Cayley graph of $C_{n}\rtimes C_{r}$.
Proof.
It is known that since $\gcd(k,n)=1$ and $k^{r}=1\mod n$ the semidirect product $C_{n}\rtimes C_{r}$ has the (minimal) presentation:
$<a,b\mid a^{r}=b^{n}=e,aba^{-1}=b^{k}>.$
We show that $G(n,k)$ is isomorphic to an induced subgraph of $\textrm{Cay}(C_{n}\rtimes C_{r},\{a,b\})$. Indeed, for $10\leq i\leq n-$ 1 map every outer vertex $u_{i}$ to $b^{i}$ and every inner vertex $v_{i}$ to $b^{i}a$. See the left side of Figure 2 for an example. Clearly, by minimality of the generating set, the outer vertices induce a cycle of length $n$ in $\textrm{Cay}(C_{n}\rtimes C_{r},\{a,b\})$. Further, there is an edge between $u_{i}$ and $v_{j}$ if and only if $b^{-i}b^{j}a\in\{a,b\}$ or $a^{-1}b^{-j}b^{i}\in\{a,b\}$. This is the case if and only if $i=j$. Now, consider inner vertices $v^{i}v^{j}$ these have an edge if $a^{-1}b^{-i}b^{j}a\in\{a,b\}$ or $a^{-1}b^{-j}b^{i}a\in\{a,b\}$. Using $aba^{-1}=b^{k}$ one gets that this is the case if and only if $i-j=k\mod n$, which finishes the proof. ∎
Note that the characterization of generalized Petersen graphs, that are minimal Cayley graphs [20, 18] yields a subset of the graphs of Theorem 5.1. Apart from the above result, we further checked for some small parameters if a given generalized Petersen graph is induced subgraph of a minimal Cayley graph of a group. For this we used the lists of such graphs computed independently by Kolja Knauer [5] and Rhys Evans [22]. We display the results in the right side of Figure 2.
6 Questions
Our work leaves two principal questions open:
Conjecture 6.1.
For every integer $g\geq 3$ there is a connected cubic graph of girth $g$ that is not a subgraph of a minimal Cayley graph.
Conjecture 6.2.
Every generalized Petersen graph is an induced subgraph of a minimal Cayley graph.
Acknowledgements:
KK was supported by the MICINN grant PID2022-137283NB-C22 and by the Spanish State Research Agency, through the Severo Ochoa and María de Maeztu Program for Centers and Units of Excellence in R&D (CEX2020-001084-M).

### v2

@@
 For a group $\Gamma$ and subset $S$ the Cayley digraph $\overrightarrow{\textrm{Cay}}(\Gamma,S)$ has vertex set $\Gamma$, where two elements $a,b\in\Gamma$ form an arc $(a,b)$ if $a^{-1}b\in S$.
-Note that antiparallell arcs, i.e., of the form $(a,b)$ and $(b,a)$, may occur.
+Note that antiparallel arcs, i.e., of the form $(a,b)$ and $(b,a)$, may occur.
 We say that $D$ is a minimal Cayley digraph if there is an inclusion-minimal generating set if sone group $\Gamma$ such that $D\cong\overrightarrow{\textrm{Cay}}(\Gamma,S)$.
-Cleary,
+Clearly,
 Observation 2.1.
@@
   1. 1.
-    for every $i\in S$ the subdigraph $f^{-1}(i)$ has maximum outdegree and maximum indegree at most $1$
+    for every $i\in S$ the subdigraph $f^{-1}(i)$ has maximum outdegree and maximum indegree at most 11
   2. 2.
@@
   2. 2.
-    all cycles in $f^{-1}(i)$ are directed (by Definition 2.1.1.) and monchromatic (by Definition 2.1.3.).
+    all cycles in $f^{-1}(i)$ are directed (by Definition 2.1.1.) and monochromatic (by Definition 2.1.3.).
 Lemma 2.3.
@@
 Now, we can apply Lemma 2.3 to $D$.
-By Observatiobn 2.2.2 we get:
+By Observation 2.2.2 we get:
 Lemma 2.5.
@@
 2Input:
-[…]
-[…]
+3cc: the color currently being extended.
+4f_validf\_valid: True if $f^{-1}(c)$ satisfies Definition 2.2, False otherwise.
 5
-6if all edges are colored and $f\_valid$ then
+6if all edges are colored and f_validf\_valid then
 7 Output current coloring $f$
@@
 11 else
-12       if $f\_valid$ then
+12 if f_validf\_valid then
     // Start coloring with a new color $c+1$
@@
 27 if $\Delta(f^{-1}(c))\leq 2$ and all cycles in $f^{-1}(c)$ have the same length then
-             // Check Observation 2.6
+28
+     1ex// Check Observation 2.6
 29    if $u^{\prime},v^{\prime}$ are in different components of $G\setminus f^{-1}(c)$ for all $e^{\prime}\in f^{-1}(c)$ then
@@
     // Reset edge if invalid
-34
+35
 return False
@@
 By Observation 2.2 in each color, one of these orientations can be arbitrarily fixed.
-Further, note that by Obervation 2.4, only if some color $f^{-1}(i)$ is a matching, then its edges can potentially be completely bi-oriented.
+Further, note that by Observation 2.4, only if some color $f^{-1}(i)$ is a matching, then its edges can potentially be completely bi-oriented.
 We initialize $D=G$ unoriented and $c=1$.
-1 Input: An integer $c$ indicating the next color class to orient.
+1Input: An integer $c$ indicating the next color class to orient.
 2
@@
 9      Orient each $H\notin S$ as $H^{-}$ in $D$
-10             extendCandidateOrientation$(c+1)$
+10    extendCandidateOrientation(c+1)(c+1)
 11 if $f^{-1}(c)$ is a matching then
 12      Bi-orient all edges in $f^{-1}(c)$ in $D$
-13             extendCandidateOrientation$(c+1)$
+13    extendCandidateOrientation(c+1)(c+1)
 Algorithm 2 extendCandidateOrientation: Extends candidate orientations of a graph with no-lonely-color $k$-coloring.
 For each of the generated candidate orientations we use the following algorithm that checks the special case of Definition 2.1.3, where only cycles are considered.
-1 Input:
-[…]
+1Input:
+2D=(V,A),fD=(V,A),f: A colored directed graph, where $f$ assigns a color to each arc.
 3 Output:
 4True if, for every cycle $C$ in $D$, the color/direction pattern of $C$ starting at any vertex yields a cycle; False otherwise.
-5for every cycle $C$ in $D$ do
+50.3cm
+6for every cycle $C$ in $D$ do
 7   Represent $C$ as a sequence of arcs with directions:
 8    $C=(a^{\sigma(1)}_{1},\ldots,a^{\sigma(\ell)}_{\ell})$
-8       where each $a_{i}$ is an arc and $\sigma(i)\in\{+1,-1\}$ indicates direction. Let $\mathbf{c}=(f(a_{1}),\ldots,f(a_{\ell}))$ be the color sequence.
-9       Let $\boldsymbol{\sigma}=(\sigma(1),\ldots,\sigma(\ell))$ be the direction sequence. for every vertex $v$ in $V$ do
+9   where each $a_{i}$ is an arc and $\sigma(i)\in\{+1,-1\}$ indicates direction.
+10  1exLet $\mathbf{c}=(f(a_{1}),\ldots,f(a_{\ell}))$ be the color sequence.
+11   Let $\boldsymbol{\sigma}=(\sigma(1),\ldots,\sigma(\ell))$ be the direction sequence.
+12  1exfor every vertex $v$ in $V$ do
 13      Try to follow a path from $v$ using color sequence $\mathbf{c}$ and direction sequence $\boldsymbol{\sigma}$.
@@
 Algorithm 3 checkCycles: Verify if the color sequence of every cycle, when followed from any vertex, yields a cycle
-4 Two cubic graphs of grith 5
-There are two cubic graphs of grith 5.
+4 Two cubic graphs of girth 5
+There are two cubic graphs of girth 5.
 Our implementation [24] applied to the triplex graph, see the left of Figure 1, shows that it is not the subgraph of a minimal Cayely graph.
@@
 Figure 1: Left: The triplex graph is not subgraph of a minimal Cayley graph. Right: An orientation of the twinplex graph with a coloring from our algorithm.
-The second cubic graphs of grith 5 is the twinplex graph depicted on the right of Figure 1.
+The second cubic graphs of girth 5 is the twinplex graph depicted on the right of Figure 1.
 In this case, our procedure finds an orientation and a coloring that passes our algorithm, as depicted in the figure.
@@
 5 Generalized Petersen Graphs
-[…]
+Let $k,n$ be integers such that $0<k<\frac{n}{2}$.
+The generalized Petersen graph $G(n,k)$ is the cubic graph on vertex set $V=V_{I}\cup V_{O}$, where $V_{I}=\{v_{0},\ldots,v_{n-1}\}$ is the set of inner vertices and $V_{O}=\{u_{0},\ldots,u_{n-1}\}$ the set of outer vertices.
+The edge set is partitioned into three parts (all subscripts are considered modulo $n$): the edges $E_{O}(n,k)=\{u_{i}u_{i+1}\,|\,0\leq i\leq n-1\}$ inducing a cycle of length $n$; the edges $E_{I}(n,k)=\{v_{i}v_{i+k}\,|\,0\leq i\leq n-1\}$, inducing $\gcd(n,k)$ cycles of length $\frac{n}{\gcd(n,k)}$; and the edges $E_{S}(n,k)=\{u_{i}v_{i}\,|\,0\leq i\leq n-1\}$ forming a perfect matching of $G(n,k)$.
+Generalized Petersen graphs were introduced by Coxeter [6] and named by Watkins [27].
+Many algebraic properties of $G(n,k)$ depend on the particular $k,n$, e.g., isomorphisms [26], automorphism groups, edge-and vertex-transitivity [9], being a core [11], and being a Cayley graph [20, 18].
 Figure 2: Left: the filled vertices induce $G(7,2)$ in $\textrm{Cay}(C_{7}\rtimes C_{3},\{a,b\})$.
@@
 It is known that since $\gcd(k,n)=1$ and $k^{r}=1\mod n$ the semidirect product $C_{n}\rtimes C_{r}$ has the (minimal) presentation:
-$<a,b\mid a^{r}=b^{n}=e,aba^{-1}=b^{k}>.$
+$a,b\mid a^{r}=b^{n}=e,aba^{-1}=b^{k}>.$
 We show that $G(n,k)$ is isomorphic to an induced subgraph of $\textrm{Cay}(C_{n}\rtimes C_{r},\{a,b\})$.
-Indeed, for $10\leq i\leq n-$ 1 map every outer vertex $u_{i}$ to $b^{i}$ and every inner vertex $v_{i}$ to $b^{i}a$.
+Indeed, for $0\leq i\leq n-1$ map every outer vertex $u_{i}$ to $b^{i}$ and every inner vertex $v_{i}$ to $b^{i}a$.
 See the left side of Figure 2 for an example.
@@
 Every generalized Petersen graph is an induced subgraph of a minimal Cayley graph.
-Acknowledgements:
+Acknowledgments:
 KK was supported by the MICINN grant PID2022-137283NB-C22 and by the Spanish State Research Agency, through the Severo Ochoa and María de Maeztu Program for Centers and Units of Excellence in R&D (CEX2020-001084-M).